OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2).

OpenStudy (anonymous):

the line can also be represented as : \[ y= \frac{2}{3}x-1 \] therefore it's slope at any point is 2/3 Now to find the slope of the line that is perpendicular to this line I use: \[m_n *\frac{2}{3}=-1\] where m_n is normal. Then you know the slope and you can match it to the given points

OpenStudy (anonymous):

what I get is: \[f(x)=-\frac{3}{2}x-10\]